Together Sophie and Clarissa can paint a house in 6 days. It takes Clarissa 5 days less than Sophie to paint the house by herself. How many days does it take Sophie to paint the house by herself?
Answer by nerdybill(7384)

Together Sophie and Clarissa can paint a house in 6 days. It takes Clarissa 5 days less than Sophie to paint the house by herself. How many days does it take Sophie to paint the house by herself?
.
Let x = time (days) it takes for Sophie to do it alone
then
x-5 = time it takes for Clarissa
.
6(1/x + 1/(x-5)) = 1
multiplying both sides by x(x-5):
6(x-5 + x) = x(x-5)
6(2x-5) = x(x-5)
12x-30 = x^2-5x
-30 = x^2-17x
0 = x^2-17x+30
0 = (x-15)(x-2)
x = {15, 2}
the 2 won't make sense, throw it out leaving:
x = 15 days (Sophie's time)
.
Clarissa:
x-5 = 15-5 = 10 days